OpenStudy (maheshmeghwal9):

Prove that for any real numbers a, b, c, & d the inequality \[\color{red}{a^4+b^4+c^4+d^4\ge 4abcd.}\]is valid.

OpenStudy (maheshmeghwal9):

I have a solution to it:)

OpenStudy (ujjwal):

Then what do you need?

OpenStudy (maheshmeghwal9):

how come he gt \[a^2b^2+c^2d^2\ge 2\sqrt{a^2b^2c^2d^2}.\]

OpenStudy (ujjwal):

\[a^2+b^2\geq 2ab\]

OpenStudy (maheshmeghwal9):

k:) i gt it:)

OpenStudy (maheshmeghwal9):

thanx:)

OpenStudy (ujjwal):

yw